OpenStudy (anonymous):

Identify the vertex for the graph of y = 3x2 + 12x + 5. (2, 41) (2, 35) (−2, −31) (−2, −7)

OpenStudy (anonymous):

use \(x=-\frac{b}{2a}=-\frac{12}{2\times 3}=-2\) for the first coordinate second coordinate is what you get when you replace \(x\) by \(-2\)

OpenStudy (anonymous):

I graphed the equation, so I got -2, -7 as the vertex. D is your answer. :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome. :)